OpenStudy (anonymous):

Find h'(3) if h=f o g, f(x)=7x^2+10 and the equation of the tangent line of g at 3 is y=−3x+4.

OpenStudy (anonymous):

\(h'(x)=f'(g(x))g'(x)\) is a start

OpenStudy (anonymous):

you are told that the equation of the tangent line of \(g\) at \(x=3\) is \(y=-3x+4\) which is a round about way of telling you that \(g'(3)=-3\)

OpenStudy (anonymous):

it also tells you that \(g(3)=-3\times 3+4=-5\)

OpenStudy (queelius):

Tangent line: y = g(3) + g'(3)(x - 3); g'(3) = -3, and we are told y = -3x + 4, so g(3) + g'(3)(-3) = 4 (4 from the tangent line eq)... so, g(3) + 9 = 4, and therefore g(3) = 4 - 9 = -5. So, f'(g(x))g'(x) = 14(-5)(-3).

OpenStudy (queelius):

Looks like satellite was beating me to the punch, actually, but I had did my work on a sheet of paper before seeing his. :p

OpenStudy (anonymous):

now \[f'(x)=14x\] and so \[f'(g(x))=14\times -5=-70\] so now we are good to go \[h'(3)=f'(g(3))g'(3)=f'(-5)\times (-3)=-70\times (-3)\] nope i got beat!

OpenStudy (queelius):

satellite showed his work much more clearly than I... well done.